Sigma Corporation applies overhead cost to jobs on the basis of direct labor cost. Job V, which was started and completed during the current period, shows charges of $6,100 for direct materials, $8,900 for direct labor, and $4,895 for overhead on its job cost sheet. Job W, which is still in process at year-end, shows charges of $3,400 for direct materials and $5,400 for direct labor.

Required:

1a. Should any overhead cost be applied to Job W at year-end?

  • Yes

  • No

1b. How much overhead cost should be applied to Job W?

2. How will the costs included in Job W’s job cost sheet be reported within Sigma Corporation’s financial statements at the end of the year?

  • Raw Materials or Work-in-Process or Finished Goods
Answer #1

1a) Yes, Any overhead cost applied to Job W

1b) Applied overhead = 5400*55% = 2970

2) Job W is work in process inventory = 3400+5400+2970 = 11770